At December 31, 2016, Flint Corporation Imports reported this information on its balance sheet.

Accounts receivable $560,400
Less: Allowance for doubtful accounts 37,640


During 2017, the company had the following transactions related to receivables.

1. Sales on account $2,968,300
2. Sales returns and allowances 92,202
3. Collections of accounts receivable 2,518,800
4. Write-offs of accounts receivable deemed uncollectible 43,310
5. Recovery of bad debts previously written off as uncollectible 15,840

Compute the accounts receivable turnover. (Round answer to 1 decimal place, e.g. 12.5.)

Accounts receivable turnover

Type your answer here

times

Compute the average collection period. (Round answer to 1 decimal place, e.g. 12.5.)

Average collection period

Type your answer here

days
Answer #1

Net credit sales = 2968300 - 92202 = $2876098

Net accounts receivable balance for 2017 = (560,400-37640)+2876098 - 2,518,800 - 43,310 = $836748

Accounts receivable turnover = Net credit sales / Average accounts receivables

= $2876098 / ((560,400-37640)+836748}/2

= 4.2 times

Average collection period = 365 / 4.2 = 86.9 days